An airplane is with the wind can cover 500 miles in 2 hours. the return trip against the wind cover 2.5 hours. what is the speed of the plane? what is the speed of the air?

Answer by lwsshak3(11628)

An airplane is with the wind can cover 500 miles in 2 hours. the return trip against the wind cover 2.5 hours. what is the speed of the plane? what is the speed of the air?
***
let x=speed of plane in still air
let c=speed of the wind
x+c=speed of plane with the wind
x-c=speed of plane against the wind
travel time*speed=distance
...
2(x+c)=500
2.5(x-c)=500
..
2x+2c=500
2.5x-2.5c=500
..
5x+5c=1250 (mult 1st eq by 2.5)
5x-5c=1000 (mult 2nd eq by 2.0)
add
10x=2250
x=225
2c=500-2x=500-450=50
c=25
..
speed of plane in still air=225 mph
speed of the wind=25 mph